# Cosmic Ray Magnetohydrodynamics: A New Two-Moment Framework with Numerical Implementation

### v1

@@ 5 Numerical Tests of the full CR-MHD system
  • The CR-modified acoustic wave can become unstable when $v_{A}>C_{s}$, consistent with previous findings by Begelman & Zweibel (1994) and Tsung et al. (2023).
@@ 6 Discussion
6.1 Links to previous two-moment frameworks
6.1.1 Comparison with Jiang & Oh (2018)
The numerical scheme in Jiang & Oh (2018) is conceived within the context of waves generated by CRSI, such that only Alfvén waves along one direction can be present (either along or opposite to the local $B$ field, but not both), with the CR pressure anisotropy $\Delta P_{\rm cr}$ not included. We show that in this special case, our equations can reduce to the two-moment scheme in Jiang & Oh (2018).
@@ 6.1.1 Comparison with Jiang & Oh (2018)
We note that Jiang & Oh (2018) use large perpendicular scattering coefficients to constrain $\vec{F}_{\rm cr}-(\mathcal{E}_{\rm cr}+P_{\rm cr})\vec{u}$ aligned with $\vec{b}$, so that the Lorentz force terms vanish.
@@
On the other hand, moving the $\vec{v}_{\rm st}$ term in equation Equation 86 to the LHS provides (note that ignoring $\Delta P_{\rm cr}$ gives $\nabla\cdot\mathbf{P}_{\rm cr}=\nabla P_{\rm cr}$):
@@
Jiang & Oh (2018) then defined an effective scattering coefficient $\sigma^{*}$ as
@@
Equation 87 and Equation 90 together with the definition Equation 89 are identical to the equation (5) and (10) in Jiang & Oh (2018), provided $\vec{F}_{\rm cr}$ is dominantly aligned with $\vec{b}$ with $\Delta P_{\rm cr}=0$. We thus can include their equations as a special case of our formulation, where the CR transport is dominated by CRSI and the effect of CR pressure anisotropy is not included.
6.1.2 Comparison with Thomas & Pfrommer (2019)
The formulation in Thomas & Pfrommer (2019) adopts a first-principles starting point similar to ours. The major difference is that, by truncating at the first-order Legendre polynomial to model drift anisotropy in the CR distribution, their approach precludes pressure anisotropy modes. It is therefore straightforward to reduce to their formulation simply by setting $\Delta P_{\rm cr}=0$ in our equations, which yields:
@@ 6.1.2 Comparison with Thomas & Pfrommer (2019)
These two equations are identical to their lab-frame CR equations (see equation E13 and E14 in Thomas & Pfrommer (2019), or equation 31 and 32 in Ruszkowski & Pfrommer (2023)).
6.2 Limitations and Future Extensions
@@ Acknowledgements
We are especially grateful to Pinghui Huang for sharing his multifluid dust module in Athena++ as a template for our implementation. We also thank Minghao Guo, Zitao Hu, and Shengtang Wang for their helpful suggestions during the development process, and Xiaochen Sun, Shuzhe Zeng, Suoqing Ji, Philipp Kempski, Chang-Goo Kim, Hanjue Zhu, and Alexander Wagner for valuable scientific discussions and suggestions. This work is supported by the National Science Foundation of China under grant No. 12325304, 12342501, and in part by grant 510940 from the Simons Foundation to ECO.
@@ Appendix A CR-MHD waves
In this appendix we calculate characteristic waves in a CR-MHD system. This topic has been studied by Begelman & Zweibel (1994); Tsung et al. (2022), which focused on unstable waves arising from CR-modified acoustic instabilities. Here we perform a separate derivation mainly serving two objectives: (1) setting the stage for our full CR-MHD system tests in Section 5, where we validate our numerical solutions against analytical eigenmode predictions; (2) determining the maximum wave speeds for the implementation in our HLLE Riemann solver.
@@
$+\frac{\partial P_{\rm cr,0}}{\partial x}\left(\delta u-\frac{v_{A,0}\delta\rho}{2\rho_{0}}\right)$
@@ A.1 Portrait of waves in a CR-MHD system
The quartic dispersion relation Equation A14 generally has four roots corresponding to four wave branches. By fixing characteristic speeds $\widetilde{C}_{s},\ \widetilde{C}_{\rm cr}$ and $\widetilde{V}_{m}$, we numerically solve for all roots and plot their dependence on $\widetilde{\sigma}$ in Figure 9. Points marked by $\star$ denote waves initialized in our numerical tests against eigenmode predictions (Section 5). For clarity in logarithmic scaling, we show absolute values of real and imaginary parts, with dashed lines indicating originally negative values.
We discuss two distinct regimes. The top panels demonstrate the real (left) and imaginary (right) parts of normalized wave frequency $\widetilde{\omega}$ in the case where $C_{s}\gg v_{A,0}$, and all imaginary parts are negative, indicating wave damping and no instability. However, when $C_{s}\ll v_{A,0}$, as in the bottom panels, one branch (red) exhibits a positive imaginary part (bottom right), signaling exponential growth of waves. The real part of this branch approaches ${\rm Re}(\omega)/k=C_{s}$ in the $\sigma v_{A,0}/k\rightarrow 0$ limit, and we thus identify it as the CR-acoustic instability described in Begelman & Zweibel (1994) and Tsung et al. (2022).
@@
Figure 9: Roots of the dispersion relation Equation A14 are plotted as functions of $\widetilde{\sigma}$, with fixed characteristic speeds (see panel titles). The left panels display the real parts, while right panels show the imaginary parts. Absolute values are displayed in logarithmic scale, with dashed/solid lines representing originally negative/positive values, respectively. Top panels illustrate the stable regime $(\widetilde{C}_{s}\equiv C_{s}/v_{A,0}\gg 1)$, where all imaginary parts are negative, indicating no wave growth. Bottom panels depict the unstable regime $(\widetilde{C}_{s}\ll 1)$, where the red branch has a positive imaginary part, signaling exponential growth. Since the real part of this branch approaches $\widetilde{C}_{s}$ in the weak scattering limit $(\widetilde{\sigma}\rightarrow 0)$, it is identified as the CR-modified acoustic wave. In the real-part panels, we also mark (with horizontal dotted lines) other characteristic speeds to indicate asymptotic wave speeds. In the imaginary panels, the damping or growth rates are observed to asymptotically scale either linearly or inversely with $\widetilde{\sigma}$, as indicated with dotted lines. We use $\star$ to mark four points corresponding to the waves injected for testing our code in the full CR-MHD system in Section 5. Specific data for these $\star$ points are listed in Table 1.
@@ A.2 Maximum wave speeds
We observe that the numerical values of $|\mathrm{Re}(\omega)|$ are nearly identical to our asymptotic estimates, showing only a minute upward deviation. For numerical robustness, we multiply the asymptotic values by a safety factor of 1.5 in the HLLE solver, except for $\widetilde{V}_{m}/\sqrt{3}$, which would otherwise exceed the speed of light. Similarly, we adjust the two transition thresholds to $\widetilde{\sigma}=\sqrt{3}/\widetilde{V}m$ and $\widetilde{\sigma}=1/(10\widetilde{C}{\rm max})$, narrowing the intermediate regime and further overestimating the maximum wave speed used in the HLLE solver.
@@ Appendix B Stability Criterion for Time Step
In this appendix, we go through the traditional von Neumann stability analysis to provide guidelines for the time step $\Delta t$. We make several simplifications in our analysis. First, since our goal is to study the numerical instability attributed to the sign-flipping feature of the scattering coefficient (Equation 68) at $P_{\rm cr}$ extrema in the case where the scattering mechanism is subject to CRSI, we remove irrelevant terms including the Lorentz force terms and the $\Delta P_{\rm cr}$ terms in the equations. Furthermore, we set the CR energy source term zero in accordance with Equation 67. Finally, we omit the $(\sigma_{+}+\sigma_{-})F_{\rm cr}$ term in the CR flux equation because it has no sign-flipping behavior and it is therefore not the cause of the problem. After all the reductions, and adopting the form for the scattering coefficient from Equation 68 we are left with the two following equations (assuming static MHD gas and $\vec{b}=\hat{x}$):
@@
$\frac{1}{V_{m}^{2}}\frac{\partial\vec{F}_{\rm cr}}{\partial t}+\nabla\cdot\mathbf{P}_{\rm cr}=\frac{4}{3}(\sigma_{+}-\sigma_{-})\mathcal{E}_{\rm cr}v_{A}=\frac{4}{3}\sigma_{0,\rm st}\frac{|\nabla\mathcal{E}_{\rm cr}|}{v_{A}\mathcal{E}_{\rm cr}}{\rm sgn}\left(-\nabla\mathcal{E}_{\rm cr}\cdot\vec{b}\right)v_{A}\mathcal{E}_{\rm cr}=-\frac{4}{3}\sigma_{0,\rm st}\nabla\mathcal{E}_{\rm cr}$ (B2)
@@
$\frac{F_{\rm cr,i}^{(n+1)}-F_{\rm cr,i}^{(n)}}{V_{m}\Delta t}+\frac{V_{m}}{6\Delta x}\left(E_{\rm cr,i+1}^{(n)}-E_{\rm cr,i-1}^{(n)}\right)-\frac{\left(F_{\rm cr,i+1}^{(n)}+F_{\rm cr,i-1}^{(n)}-2F_{\rm cr,i}^{(n)}\right)}{2\sqrt{3}\Delta x}=-\frac{4}{3}\sigma_{0,\rm st}\frac{V_{m}}{2\Delta x}\left(E_{\rm cr,i+1}^{(n)}-E_{\rm cr,i-1}^{(n)}\right)$ (B4)
@@
$A_{k}=1-\frac{2\sqrt{3}V_{m}\Delta t}{3\Delta x}{\rm sin}^{2}\frac{k\Delta x}{2}\pm i\sqrt{\frac{1+4\sigma_{0,\rm st}}{3}}\frac{V_{m}\Delta t}{\Delta x}{\rm sin}k\Delta x$ (B5)
@@
$\frac{4\sqrt{3}}{3}\frac{V_{m}\Delta t}{\Delta x}{\rm sin}^{2}\frac{k\Delta x}{2}+\frac{4}{3}\frac{V_{m}^{2}\Delta t^{2}}{\Delta x^{2}}{\rm sin}^{4}\frac{k\Delta x}{2}+\left(\frac{1+4\sigma_{0,\rm st}}{3}\right)\frac{V_{m}^{2}\Delta t^{2}}{\Delta x^{2}}{\rm sin}^{2}k\Delta x<0$ (B6)
@@
$\frac{4}{3}\frac{V_{m}\Delta t}{\Delta x}{\rm sin}^{2}\frac{k\Delta x}{2}\left[\frac{V_{m}\Delta t}{\Delta x}+4\sigma_{0,\rm st}\frac{V_{m}\Delta t}{\Delta x}{\rm cos}^{2}\frac{k\Delta x}{2}-\sqrt{3}\right]$ (B7)
@@
$\Delta t<\frac{\sqrt{3}\Delta x}{(1+4\sigma_{0,\rm st})V_{m}}$ (B8)
@@ Appendix C CR energy evolution in cylindrical coordinates with scattering due to CRPAI
In this appendix, we provide a derivation for the $\mathcal{E}_{\rm cr}$ evolution in a cylindrical coordinate, where the CR scattering is subject to the CRPAI as described in Section 4.6. Specifically, we are trying to solve the two-moment equations:
$\frac{\partial\mathcal{E}_{\rm cr}}{\partial t}+\frac{1}{R}\frac{\partial(RF_{\rm cr})}{\partial R}=\left(\frac{4\sigma_{0,\rm aniso}-1}{3}\right)\frac{1}{B}\left|\frac{dB}{dR}\right|\mathcal{E}_{\rm cr}v_{A}$ (C1)
$\frac{1}{V_{m}^{2}}\frac{\partial F_{\rm cr}}{\partial t}+\frac{1}{3}\frac{\partial\mathcal{E}_{\rm cr}}{\partial R}=-\frac{\sigma_{0,\rm aniso}}{v_{A}}\frac{1}{B}\left|\frac{dB}{dR}\right|F_{\rm cr}$ (C2)
where $\sigma_{0,\rm aniso}$ is a positive constant, $B(R)=B_{0}(R_{0}/R)$ and $v_{A}(R)=v_{A,0}(R_{0}/R)$ with $B_{0}$, $v_{A0}$ and $R_{0}$ being constants simply for normalizations.
@@
$F_{\rm cr}=-\frac{v_{A0}R_{0}}{3\sigma_{0,\rm aniso}}\frac{\partial\mathcal{E}_{\rm cr}}{\partial R}$ (C3)
@@
$\frac{\partial\mathcal{E}_{\rm cr}}{\partial t}-\frac{v_{A0}R_{0}}{3\sigma_{0,\rm aniso}}\frac{\partial^{2}\mathcal{E}_{\rm cr}}{\partial R^{2}}-\frac{v_{A0}R_{0}}{3\sigma_{0,\rm aniso}R}\frac{\partial\mathcal{E}_{\rm cr}}{\partial R}=\left(\frac{4\sigma_{0,\rm aniso}-1}{3}\right)\frac{v_{A0}R_{0}}{R^{2}}\mathcal{E}_{\rm cr}$ (C4)
@@
$\frac{1}{T(t)}\frac{dT}{dt}=\frac{v_{A0}R_{0}}{3\sigma_{0,\rm aniso}}\frac{1}{L(R)}\left[\frac{d^{2}L}{dR^{2}}+\frac{1}{R}\frac{dL}{dR}+\left(\frac{4\sigma_{0,\rm aniso}-1}{3R^{2}}\right)\sigma_{0,\rm aniso}L\right]$ (C5)
@@
$R^{2}\frac{d^{2}L}{dR^{2}}+R\frac{dL}{dR}+\left(4\sigma_{0,\rm aniso}^{2}-\sigma_{0,\rm aniso}+\frac{3\sigma_{0,\rm aniso}}{v_{A0}R_{0}}\lambda R^{2}\right)L=0$ (C6)
@@
$L(R)=C_{2}J_{\alpha}\left(\sqrt{\frac{3\lambda\sigma_{0,\rm aniso}}{v_{A0}R_{0}}}R\right)+C_{3}Y_{\alpha}\left(\sqrt{\frac{3\lambda\sigma_{0,\rm aniso}}{v_{A0}R_{0}}}R\right)$ (C7)
where $J_{\alpha}$ and $Y_{\alpha}$ are respectively the first and second kind of Bessel functions. $\alpha\equiv\sqrt{(1-4\sigma_{0,\rm aniso})\sigma_{0,\rm aniso}}$ is the order, and $C_{1}$, $C_{2}$ and $C_{3}$ are integral constants.
In Section 4.6, for simplicity we choose $C_{3}=0$, $\sigma_{0,\rm aniso}=0.25$, and $v_{A0}=R_{0}=\lambda=C_{1}=C_{2}=1$, so that one specific solution for $\mathcal{E}_{\rm cr}$ is $\mathcal{E}_{\rm cr}(t,R)=T(t)L(R)=J_{0}(\sqrt{3}R/2)e^{-t}$, which we send in as initial condition to test the performance our code.

### v2

@@ V Numerical Tests of the full CR-MHD system
  • The CR-modified acoustic wave can become unstable when $v_{A}>C_{s}$, consistent with previous findings by Begelman and Zweibel (1994) and Tsung et al. (2023).
@@ VI Discussion
VI.1 Links to Previous Two-Moment Frameworks
VI.1.1 Comparison with Jiang and Oh (2018)
The numerical scheme in Jiang and Oh (2018) is conceived within the context of waves generated by CRSI, such that only Alfvén waves along one direction can be present (either along or opposite to the local $B$ field, but not both), with the CR pressure anisotropy $\Delta P_{\rm cr}$ not included. We show that in this special case, our equations can reduce to the two-moment scheme in Jiang and Oh (2018).
@@ VI.1.1 Comparison with Jiang and Oh (2018)
We note that Jiang and Oh (2018) use large perpendicular scattering coefficients to constrain $\vec{F}_{\rm cr}-(\mathcal{E}_{\rm cr}+P_{\rm cr})\vec{u}$ aligned with $\vec{b}$, so that the Lorentz force terms vanish.
@@
On the other hand, moving the $\vec{v}_{\rm st}$ term in Equation 87 to the LHS provides (note that ignoring $\Delta P_{\rm cr}$ gives $\nabla\cdot\mathbf{P}_{\rm cr}=\nabla P_{\rm cr}$):
@@
Jiang and Oh (2018) then defined an effective scattering coefficient $\sigma^{*}$ as
@@
Equation 88 and Equation 91 together with the definition Equation 90 are identical to the equation (5) and (10) in Jiang and Oh (2018), provided $\vec{F}_{\rm cr}$ is dominantly aligned with $\vec{b}$ with $\Delta P_{\rm cr}=0$. We thus can include their equations as a special case of our formulation, where the CR transport is dominated by CRSI and the effect of CR pressure anisotropy is not included.
VI.1.2 Comparison with Thomas and Pfrommer (2019)
The formulation in Thomas and Pfrommer (2019) adopts a first-principles starting point similar to ours. One key formal difference is that, by truncating at the first-order Legendre polynomial to model drift anisotropy in the CR distribution, their approach precludes CR pressure anisotropy modes and the associated microphysics. It is therefore straightforward to reduce to their formulation simply by setting $\Delta P_{\rm cr}=0$ in our equations, which yields:
@@ VI.1.2 Comparison with Thomas and Pfrommer (2019)
These two equations are identical to their lab-frame CR equations (see equation E13 and E14 in Thomas and Pfrommer (2019), or equation 31 and 32 in Ruszkowski and Pfrommer (2023)).
Another important conceptual distinction lies in how scattering coefficients $\sigma_{\pm}^{L,R}$ are determined. Thomas and Pfrommer (2019) co-evolve the Alfvén wave subsystem dynamically alongside CR-MHD equations to obtain instantaneous $\sigma_{\pm}^{L,R}$ associated with Alfvén wave intensities, rather than prescribing a local steady-state balance as in our work. Their approach thus naturally accommodates regimes such as “Alfvén-wave dark regions” (Thomas et al., 2023), where $\nabla P_{\rm cr}$ is nearly perpendicular to the magnetic field, and the assumption of rapid wave growth/damping balance no longer holds due to suppressed gyro-resonant instabilities.
In typical ISM conditions, however, wave growth from gyro-resonant instabilities (e.g., CRSI and CRPAI) saturates on timescales much shorter than the dynamical timescale resolved in large-scale simulations, with the wave subsystem converging toward QLT predictions. More subtly, our steady-state scattering coefficients can be directly calibrated against MHD-PIC simulations (e.g., Bai, 2022; Sun et al., 2024), capturing deviations from QLT predictions. By comparison, calibrating the dynamically evolving wave subsystem is less straightforward. We therefore view the two methods as complementary, with the steady-state framework providing a practical and robust tool for astrophysical applications where rapid kinetic equilibration is expected.
VI.2 Limitations and Future Work
@@
In this work, we focus on idealized tests without introducing physical units. These tests are designed to isolate physical effects and verify the numerical implementation and performance of our CR module, as demonstrated in this work. Future efforts will apply this framework to realistic astrophysical environments, with such applications presented in forthcoming work.
@@ Acknowledgements
We are especially grateful to Pinghui Huang for sharing his multifluid dust module in Athena++ as a template for our implementation. We also thank Minghao Guo, Zitao Hu, and Shengtang Wang for their helpful suggestions during the development process, and Phil Hopkins, Xiaochen Sun, Shuzhe Zeng, Suoqing Ji, Philipp Kempski, Chang-Goo Kim, Hanjue Zhu, and Alexander Wagner for valuable scientific discussions and suggestions. We thank the anonymous referee for thoughtful and constructive comments that helped improve the manuscript.
This work is supported by the National Science Foundation of China under grant No. 12325304, 12342501, and in part by grant 510940 from the Simons Foundation to ECO.
@@ Appendix A CR-MHD waves
In this appendix we calculate characteristic waves in a CR-MHD system. This topic has been studied by Begelman and Zweibel (1994); Tsung et al. (2022), which focused on unstable waves arising from CR-modified acoustic instabilities. Here we perform a separate derivation mainly serving two objectives: (1) setting the stage for our full CR-MHD system tests in Section V, where we validate our numerical solutions against analytical eigenmode predictions; (2) determining the maximum wave speeds for the implementation in our HLLE Riemann solver.
@@ A.1 Portrait of waves in a CR-MHD system
The quartic dispersion relation Equation A14 generally has four roots corresponding to four wave branches. By fixing characteristic speeds $\widetilde{C}_{s},\ \widetilde{C}_{\rm cr}$ and $\widetilde{V}_{m}$, we numerically solve for all roots and plot their dependence on $\widetilde{\sigma}$ in Figure 9. Points marked by $\star$ denote waves initialized in our numerical tests against eigenmode predictions (Section V). For clarity in logarithmic scaling, we show absolute values of real and imaginary parts, with dashed lines indicating originally negative values. A similar dispersion analysis was also carried out by Thomas et al. (2021) for code verification.
We discuss two distinct regimes. The top panels demonstrate the real (left) and imaginary (right) parts of normalized wave frequency $\widetilde{\omega}$ in the case where $C_{s}\gg v_{A,0}$, and all imaginary parts are negative, indicating wave damping and no instability. However, when $C_{s}\ll v_{A,0}$, as in the bottom panels, one branch (red) exhibits a positive imaginary part (bottom right), signaling exponential growth of waves. The real part of this branch approaches ${\rm Re}(\omega)/k=C_{s}$ in the $\sigma v_{A,0}/k\rightarrow 0$ limit, and we thus identify it as the CR-acoustic instability described in Begelman and Zweibel (1994) and Tsung et al. (2022).
@@
Figure 9: Roots of the dispersion relation Equation A14 are plotted as functions of $\widetilde{\sigma}$, with fixed characteristic speeds (see panel titles). The left panels display the real parts, while right panels show the imaginary parts. Absolute values are displayed in logarithmic scale, with dashed/solid lines representing originally negative/positive values, respectively. Top panels illustrate the stable regime $(\widetilde{C}_{s}\equiv C_{s}/v_{A,0}\gg 1)$, where all imaginary parts are negative, indicating no wave growth. Bottom panels depict the unstable regime $(\widetilde{C}_{s}\ll 1)$, where the red branch has a positive imaginary part, signaling exponential growth. Since the real part of this branch approaches $\widetilde{C}_{s}$ in the weak scattering limit $(\widetilde{\sigma}\rightarrow 0)$, it is identified as the CR-modified acoustic wave. In the real-part panels, we also mark (with horizontal dotted lines) other characteristic speeds to indicate asymptotic wave speeds. In the imaginary panels, the damping or growth rates are observed to asymptotically scale either linearly or inversely with $\widetilde{\sigma}$, as indicated with dotted lines. We use $\star$ to mark four points corresponding to the waves injected for testing our code in the full CR-MHD system in Section V. Specific data for these $\star$ points are listed in Table 1.
@@ A.2 Maximum wave speeds
We observe that the numerical values of $|\mathrm{Re}(\omega)|$ are nearly identical to our asymptotic estimates, showing only a minute upward deviation. For numerical robustness, we multiply the asymptotic values by a safety factor of 2 in the HLLE solver, except for $\widetilde{V}_{m}/\sqrt{3}$, which would otherwise exceed the speed of light. Similarly, we adjust the two transition thresholds to $\widetilde{\sigma}=\sqrt{3}/\widetilde{V}_{m}$ and $\widetilde{\sigma}=1/(10\widetilde{C}_{\rm max})$, narrowing the intermediate regime and further overestimating the maximum wave speed used in the HLLE solver.
@@ Appendix B Stability Criterion for Time Step
In this appendix, we go through the traditional von Neumann stability analysis to provide guidelines for the time step $\Delta t$. We make several simplifications in our analysis. First, since our goal is to study the numerical instability attributed to the sign-flipping feature of the scattering coefficient (Equation 68) at $P_{\rm cr}$ extrema under CRSI-driven scattering, we remove irrelevant terms including the Lorentz force terms and the $\Delta P_{\rm cr}$ terms in the equations. Furthermore, we set the CR energy source term zero in accordance with Equation 67. Finally, we omit the $(\sigma_{+}+\sigma_{-})F_{\rm cr}$ term in the CR flux equation because it has no sign-flipping behavior and it is therefore not the cause of the problem. After all the reductions, and adopting the form for the scattering coefficient from Equation 68 we are left with the two following equations (assuming static MHD gas and $\vec{b}=\hat{x}$):
@@
$\frac{1}{V_{m}^{2}}\frac{\partial\vec{F}_{\rm cr}}{\partial t}+\nabla\cdot\mathbf{P}_{\rm cr}=\frac{4}{3}(\sigma_{+}-\sigma_{-})\mathcal{E}_{\rm cr}v_{A}=\frac{4}{3}\widetilde{\sigma}_{0,\rm st}\frac{|\nabla\mathcal{E}_{\rm cr}|}{v_{A}\mathcal{E}_{\rm cr}}{\rm sgn}\left(-\nabla\mathcal{E}_{\rm cr}\cdot\vec{b}\right)v_{A}\mathcal{E}_{\rm cr}=-\frac{4}{3}\widetilde{\sigma}_{0,\rm st}\nabla\mathcal{E}_{\rm cr}$ (B2)
@@
$\frac{F_{\rm cr,i}^{(n+1)}-F_{\rm cr,i}^{(n)}}{V_{m}\Delta t}+\frac{V_{m}}{6\Delta x}\left(E_{\rm cr,i+1}^{(n)}-E_{\rm cr,i-1}^{(n)}\right)-\frac{\left(F_{\rm cr,i+1}^{(n)}+F_{\rm cr,i-1}^{(n)}-2F_{\rm cr,i}^{(n)}\right)}{2\sqrt{3}\Delta x}=-\frac{4}{3}\widetilde{\sigma}_{0,\rm st}\frac{V_{m}}{2\Delta x}\left(E_{\rm cr,i+1}^{(n)}-E_{\rm cr,i-1}^{(n)}\right)$ (B4)
@@
$A_{k}=1-\frac{2\sqrt{3}V_{m}\Delta t}{3\Delta x}{\rm sin}^{2}\frac{k\Delta x}{2}\pm i\sqrt{\frac{1+4\widetilde{\sigma}_{0,\rm st}}{3}}\frac{V_{m}\Delta t}{\Delta x}{\rm sin}k\Delta x$ (B5)
@@
$\frac{4\sqrt{3}}{3}\frac{V_{m}\Delta t}{\Delta x}{\rm sin}^{2}\frac{k\Delta x}{2}+\frac{4}{3}\frac{V_{m}^{2}\Delta t^{2}}{\Delta x^{2}}{\rm sin}^{4}\frac{k\Delta x}{2}+\left(\frac{1+4\widetilde{\sigma}_{0,\rm st}}{3}\right)\frac{V_{m}^{2}\Delta t^{2}}{\Delta x^{2}}{\rm sin}^{2}k\Delta x<0$ (B6)
@@
$\frac{4}{3}\frac{V_{m}\Delta t}{\Delta x}{\rm sin}^{2}\frac{k\Delta x}{2}\left[\frac{V_{m}\Delta t}{\Delta x}+4\widetilde{\sigma}_{0,\rm st}\frac{V_{m}\Delta t}{\Delta x}{\rm cos}^{2}\frac{k\Delta x}{2}-\sqrt{3}\right]$ (B7)
@@
$\Delta t<\frac{\sqrt{3}\Delta x}{(1+4\widetilde{\sigma}_{0,\rm st})V_{m}}$ (B8)
@@ Appendix C CR energy evolution in cylindrical coordinates with scattering due to CRPAI
In this appendix, we provide a derivation for the $\mathcal{E}_{\rm cr}$ evolution in a cylindrical coordinate, where the CR scattering is subject to the CRPAI as described in Section IV.6. Specifically, we are trying to solve the two-moment equations:
$\frac{\partial\mathcal{E}_{\rm cr}}{\partial t}+\frac{1}{R}\frac{\partial(RF_{\rm cr})}{\partial R}=\left(\frac{4\widetilde{\sigma}_{0,\rm aniso}-1}{3}\right)\frac{1}{B}\left|\frac{dB}{dR}\right|\mathcal{E}_{\rm cr}v_{A}$ (C1)
$\frac{1}{V_{m}^{2}}\frac{\partial F_{\rm cr}}{\partial t}+\frac{1}{3}\frac{\partial\mathcal{E}_{\rm cr}}{\partial R}=-\frac{\widetilde{\sigma}_{0,\rm aniso}}{v_{A}}\frac{1}{B}\left|\frac{dB}{dR}\right|F_{\rm cr}$ (C2)
where $\widetilde{\sigma}_{0,\rm aniso}$ is a positive constant, $B(R)=B_{0}(R_{0}/R)$ and $v_{A}(R)=v_{A,0}(R_{0}/R)$ with $B_{0}$, $v_{A0}$ and $R_{0}$ being constants simply for normalizations.
@@
$F_{\rm cr}=-\frac{v_{A0}R_{0}}{3\widetilde{\sigma}_{0,\rm aniso}}\frac{\partial\mathcal{E}_{\rm cr}}{\partial R}$ (C3)
@@
$\frac{\partial\mathcal{E}_{\rm cr}}{\partial t}-\frac{v_{A0}R_{0}}{3\widetilde{\sigma}_{0,\rm aniso}}\frac{\partial^{2}\mathcal{E}_{\rm cr}}{\partial R^{2}}-\frac{v_{A0}R_{0}}{3\widetilde{\sigma}_{0,\rm aniso}R}\frac{\partial\mathcal{E}_{\rm cr}}{\partial R}=\left(\frac{4\widetilde{\sigma}_{0,\rm aniso}-1}{3}\right)\frac{v_{A0}R_{0}}{R^{2}}\mathcal{E}_{\rm cr}$ (C4)
@@
$\frac{1}{T(t)}\frac{dT}{dt}=\frac{v_{A0}R_{0}}{3\widetilde{\sigma}_{0,\rm aniso}}\frac{1}{L(R)}\left[\frac{d^{2}L}{dR^{2}}+\frac{1}{R}\frac{dL}{dR}+\left(\frac{4\widetilde{\sigma}_{0,\rm aniso}-1}{3R^{2}}\right)\widetilde{\sigma}_{0,\rm aniso}L\right]$ (C5)
@@
$R^{2}\frac{d^{2}L}{dR^{2}}+R\frac{dL}{dR}+\left(4\widetilde{\sigma}_{0,\rm aniso}^{2}-\widetilde{\sigma}_{0,\rm aniso}+\frac{3\widetilde{\sigma}_{0,\rm aniso}}{v_{A0}R_{0}}\lambda R^{2}\right)L=0$ (C6)
@@
$L(R)=C_{2}J_{\alpha}\left(\sqrt{\frac{3\lambda\widetilde{\sigma}_{0,\rm aniso}}{v_{A0}R_{0}}}R\right)+C_{3}Y_{\alpha}\left(\sqrt{\frac{3\lambda\widetilde{\sigma}_{0,\rm aniso}}{v_{A0}R_{0}}}R\right)$ (C7)
where $J_{\alpha}$ and $Y_{\alpha}$ are respectively the first and second kind of Bessel functions. $\alpha\equiv\sqrt{(1-4\widetilde{\sigma}_{0,\rm aniso})\widetilde{\sigma}_{0,\rm aniso}}$ is the order, and $C_{1}$, $C_{2}$ and $C_{3}$ are integral constants.
In Section IV.6, for simplicity we choose $C_{3}=0$, $\widetilde{\sigma}_{0,\rm aniso}=0.25$, and $v_{A0}=R_{0}=\lambda=C_{1}=C_{2}=1$, so that one specific solution for $\mathcal{E}_{\rm cr}$ is $\mathcal{E}_{\rm cr}(t,R)=T(t)L(R)=J_{0}(\sqrt{3}R/2)e^{-t}$, which we send in as initial condition to test the performance our code.